# Spin Theory Based on the Extended Least Action Principle and Information Metrics: Quantization, Entanglement, and Bell Test With Time Delay

### v2

@@ Appendix C Proof of (36) in Standard Quantum Mechanics
This state can be obtained by rotating the state of spin-up along the $z$-axis by an angle $\beta_{1}$. The rotation operator is $\hat{R}(\beta_{1})=e^{-i\beta_{1}\sigma_{z}/2}$ where $\sigma_{z}$ is the $z$-component of the Pauli operator. Suppose the electron is in this initial state, and passes a Stern-Gerlach apparatus that is configured with a magnetic field with direction tilt an angle $\beta_{2}$ from the $z$-axis. The measurement operator for spin-up along the $\beta_{2}$ direction is $\hat{O}_{+}=|\psi_{\beta_{2}}^{+}\rangle\langle\psi_{\beta_{2}}^{+}|$ where
@@ Appendix D Proof of (72)
First we consider initial condition described by the joint probability density functions are given by (63). This initial condition can be further decomposed such that half of the chance that the joint probability density $\tilde{p}_{z,1}(\theta_{A}^{z},\theta_{B}^{z})=\delta(\theta_{A}^{z})\delta(\theta_{B}^{z}-\pi)$ and the other half chance that $\tilde{p}_{z,2}(\theta_{A}^{z},\theta_{B}^{z})=\delta(\theta_{A}^{z}-\pi)\delta(\theta_{B}^{z})$. Now for the initial condition that $\tilde{p}_{z,1}(\theta_{A}^{z},\theta_{B}^{z})=\delta(\theta_{A}^{z})\delta(\theta_{B}^{z}-\pi)$, suppose Alice measures $A$ along a direction determined by unit vector $a$ which forms an angle with $z$-axis by $\alpha$, and Bob measures $B$ along a direction determined by unit vector $b$ which forms an angle with $z$-axis by $\beta$. In typical Bell experiments, the measurements of Alice and Bob are performed almost at the same time. Therefore, we will not consider the time dependency of measurement outcomes as discussed in Section V.1.
@@
The initial condition can also be described by joint probability density $\tilde{p}_{z,2}(\theta_{A}^{z},\theta_{B}^{z})=\delta(\theta_{A}^{z}-\pi)\delta(\theta_{B}^{z})$. Performing the same calculation with this initial condition, one will find the resulting $E(a,b|\tilde{p}_{z})$ is the same. Thus, the statistical average of $E(a,b|\tilde{p}_{z})$ for both initial conditions is still given by (106).
Next we consider the initial condition described by the joint probability density functions are given by (64). This initial condition can be further decomposed such that half of the chance that the joint probability density $\tilde{p}_{y,1}(\theta_{A}^{y},\theta_{B}^{y})=\delta(\theta_{A}^{y})\delta(\theta_{B}^{y}-\pi)$ and the other half chance that $\tilde{p}_{y,2}(\theta_{A}^{y},\theta_{B}^{y})=\delta(\theta_{A}^{y}-\pi)\delta(\theta_{B}^{y})$. Now for the initial condition that $\tilde{p}_{y,1}(\theta_{A}^{y},\theta_{B}^{y})=\delta(\theta_{A}^{y})\delta(\theta_{B}^{y}-\pi)$, Alice’s measurement direction of $a$ forms an angle with the $y$-axis that is given by $\alpha^{\prime}=\pi/2-\alpha$. Similarly, Bob’s measurement direction of $b$ forms an angle with the $y$-axis that is given by $\beta^{\prime}=\pi/2-\beta$. The calculation for $E(a,b|y)$ is exactly the same as calculation for $E(a,b|z)$ but with the replacement of $\alpha\to\alpha^{\prime}$ and $\beta\to\beta^{\prime}$. Thus,
@@
Similarly, the initial condition can also be described by joint probability density $\tilde{p}_{y,2}(\theta_{A}^{y},\theta_{B}^{y})=\delta(\theta_{A}^{y}-\pi)\delta(\theta_{B}^{y})$, but the calculation result for $E(a,b|\tilde{p}_{y})$ is the same. Thus, the statistical average of $E(a,b|\tilde{p}_{y})$ for both initial conditions is still given by (107).
@@
Now consider the situation that Bob delays his spin measurement on electron $B$ by a time period $\Delta t$. When $\Delta t$ is sufficiently large, even though Alice performs measurement on $A$ and obtains spin-up, Bob’s measurement after $\Delta t$ can obtain spin-down or spin-up because the initial orientation for $B$ can be either $\theta_{B}^{z}\in\Theta^{+}_{z}$ or $\theta_{B}^{z}\in\Theta^{-}_{z}$. Thus, Bob’s measurement outcome can be
@@
Therefore, the expectation value $E(a,b|\tilde{p}_{z})$ can be $E(a,b|\tilde{p}_{z})=-\cos(\alpha)\cos(\beta)$, or $E(a,b|\tilde{p}_{z})=\cos(\alpha)\cos(\beta)$ with equal probability. This resulting the average $E(a,b|\tilde{p}_{z})=0$. Suppose the correlation between $\theta_{A}^{y}$ and $\theta^{y}_{B}$ still holds, so that (107) is still valid. The overall expectation value becomes
@@
If the correlation between $\theta_{A}^{y}$ and $\theta^{y}_{B}$ does not hold either after the delay $\Delta t$, and $\theta_{B}^{y}\in\Theta^{+}_{y}$ or $\theta_{B}^{y}\in\Theta^{-}_{y}$ randomly, then one can calculate that $E(a,b|\tilde{p}_{y})=0$. This results in $E(a,b|\tilde{p})=0$ regardless of the configuration of $a$ and $b$. Clearly, in such a scenario, the Bell-CHSH inequality is not violated.

### v3

@@ Appendix D Proof of (36) in Standard Quantum Mechanics
This state can be obtained by rotating the state of spin-up along the $z$-axis by an angle $\beta_{1}$. The rotation operator is $\hat{R}(\beta_{1})=e^{-i\beta_{1}\sigma_{z}/2}$ where $\sigma_{z}$ is the $z$ component of the Pauli operator. Suppose the electron is in this initial state, and passes a Stern-Gerlach apparatus that is configured with a magnetic field with direction tilt an angle $\beta_{2}$ from the $z$-axis. The measurement operator for spin-up along the $\beta_{2}$ direction is $\hat{O}_{+}=|\psi_{\beta_{2}}^{+}\rangle\langle\psi_{\beta_{2}}^{+}|$ where
@@ Appendix E Proof of (71)
First we consider the initial condition described by the joint probability density functions given by (61). This initial condition can be further decomposed such that half of the chance that the joint probability density $\tilde{p}_{z,1}(\theta_{A}^{z},\theta_{B}^{z})=\delta(\theta_{A}^{z})\delta(\theta_{B}^{z}-\pi)$ and the other half chance that $\tilde{p}_{z,2}(\theta_{A}^{z},\theta_{B}^{z})=\delta(\theta_{A}^{z}-\pi)\delta(\theta_{B}^{z})$. Now for the initial condition $\tilde{p}_{z,1}(\theta_{A}^{z},\theta_{B}^{z})=\delta(\theta_{A}^{z})\delta(\theta_{B}^{z}-\pi)$, suppose that Alice measures $A$ along a direction determined by unit vector $a$ which forms an angle with the $z$-axis by $\alpha$, and Bob measures $B$ along a direction determined by unit vector $b$ which forms an angle with $z$-axis by $\beta$. In typical Bell experiments, the measurements of Alice and Bob are performed almost at the same time. Therefore, we will not consider the time dependency of measurement outcomes as discussed in Section V.1.
@@
The initial condition can also be described by the joint probability density $\tilde{p}_{z,2}(\theta_{A}^{z},\theta_{B}^{z})=\delta(\theta_{A}^{z}-\pi)\delta(\theta_{B}^{z})$. Performing the same calculation with this initial condition, one will find that the resulting $E(a,b|\tilde{p}_{z})$ is the same. Thus, the statistical average of $E(a,b|\tilde{p}_{z})$ for both initial conditions is still given by (116).
Next, we consider the initial condition described by the joint probability density functions given by (62). This initial condition can be further decomposed such that half of the chance that the joint probability density $\tilde{p}_{y,1}(\theta_{A}^{y},\theta_{B}^{y})=\delta(\theta_{A}^{y})\delta(\theta_{B}^{y}-\pi)$ and the other half chance that $\tilde{p}_{y,2}(\theta_{A}^{y},\theta_{B}^{y})=\delta(\theta_{A}^{y}-\pi)\delta(\theta_{B}^{y})$. Now for the initial condition that $\tilde{p}_{y,1}(\theta_{A}^{y},\theta_{B}^{y})=\delta(\theta_{A}^{y})\delta(\theta_{B}^{y}-\pi)$, Alice’s measurement direction of $a$ forms an angle with the $y$-axis given by $\alpha^{\prime}=\pi/2-\alpha$. Similarly, Bob’s measurement direction of $b$ forms an angle with the $y$-axis given by $\beta^{\prime}=\pi/2-\beta$. The calculation for $E(a,b|y)$ is exactly the same as the calculation for $E(a,b|z)$ but with the replacement of $\alpha\to\alpha^{\prime}$ and $\beta\to\beta^{\prime}$. Thus,
@@
Similarly, the initial condition can also be described by the joint probability density $\tilde{p}_{y,2}(\theta_{A}^{y},\theta_{B}^{y})=\delta(\theta_{A}^{y}-\pi)\delta(\theta_{B}^{y})$, but the calculation result for $E(a,b|\tilde{p}_{y})$ is the same. Thus, the statistical average of $E(a,b|\tilde{p}_{y})$ for both initial conditions is still given by (117).
@@
Now consider the situation in which Bob delays his spin measurement on electron $B$ by a time period $\Delta t$. When $\Delta t$ is sufficiently large, even though Alice performs the measurement on $A$ and gets spin-up, Bob’s measurement after $\Delta t$ can obtain spin-down or spin-up because the initial orientation for $B$ can be $\theta_{B}^{z}\in\Theta^{+}_{z}$ or $\theta_{B}^{z}\in\Theta^{-}_{z}$. Thus, Bob’s measurement outcome can be
@@
Therefore, the expectation value $E(a,b|\tilde{p}_{z})$ can be $E(a,b|\tilde{p}_{z})=-\cos(\alpha)\cos(\beta)$, or $E(a,b|\tilde{p}_{z})=\cos(\alpha)\cos(\beta)$ with the same probability. This results in the average $E(a,b|\tilde{p}_{z})=0$. Suppose that the correlation between $\theta_{A}^{y}$ and $\theta^{y}_{B}$ still holds, so that (117) is still valid. The overall expectation value becomes
@@
If the correlation between $\theta_{A}^{y}$ and $\theta^{y}_{B}$ does not hold either after delay $\Delta t$, and $\theta_{B}^{y}\in\Theta^{+}_{y}$ or $\theta_{B}^{y}\in\Theta^{-}_{y}$ randomly, then one can calculate $E(a,b|\tilde{p}_{y})=0$. This results in $E(a,b|\tilde{p})=0$ regardless of the configuration of $a$ and $b$. Clearly, in such a scenario, the Bell-CHSH inequality is not violated.